OpenStudy (anonymous):

Determine the solution set of (x - 7)2 - 144 = 0

OpenStudy (lollylau):

x-7=72 x=65

OpenStudy (lollylau):

or was it raised to a power

OpenStudy (mertsj):

Is it (x-7) squared?

OpenStudy (lollylau):

if yes, then x^2-14x-95=0

OpenStudy (anonymous):

yes it is squared

OpenStudy (anonymous):

{-5, 19} {5, -19} {5, 19} these are the answers

OpenStudy (mertsj):

Then x-7 = positive 12 or negative 12 So x = 19 or -5

OpenStudy (mertsj):